OpenStudy (anonymous):

Gene has a gasoline budget of $300 per month. He uses an average of $5 of gasoline each day he drives. Which of the following equations represents how much money is left in his gasoline budget after x days of driving?

OpenStudy (danjs):

starts with 300 take away the amount of gas he uses

OpenStudy (danjs):

The amount of gas is 5 dollars per day, 5x , where x is the number of days

OpenStudy (danjs):

what does that leave you with?

OpenStudy (danjs):

Amount left in budget = ?

OpenStudy (anonymous):

A. y = 300 - 5x B. y = 300 + 5x C. y = 300x - 5 D. y = 5x - 300

OpenStudy (danjs):

Which one starts with 300 and subtracts 5 dollars a day for x # of days?

OpenStudy (danjs):

5 dollars a day for x days = 5 times x

OpenStudy (danjs):

y = (starting amount) - 5*(Days of driving, x)

OpenStudy (anonymous):

so A ?

OpenStudy (danjs):

yes

OpenStudy (danjs):

To test if your starting amount is correct, let x = number of days = zero, you should get the starting amount of 300

OpenStudy (danjs):

At day zero, you have 300, the starting amount